Find the limit. ifit exists. 1-16 limz-4
Answer #1

0²-16 L = lim 274 2-4 lim = ²-4² 26+4 X-4 =lim (x-4) (x+4) (2-4) HA 4 /x+4) lim 144 = 4 + 4 ha 8